the roses cost $15 per dozen and the carnations cost $8 per dozen. shereda bought a total of 17 dozen flowers and paid a total of $192. how many roses did she buy?
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16949)

the roses cost $15 per dozen and the carnations cost $8 per dozen. shereda bought a total of 17 dozen flowers and paid a total of $192. how many roses did she buy?
..
Roses $15 per dozen
carnations $8 per dozen
..
total 17 dozens - $192
..
assume she bought x doz. roses
so she bought 17-x doz. carnations
..
cost of roses + cost of carnations = 192
15x+8*(17-x)=192
15x+136-8x=192
7x+136=192
add -132
7x+136-136=192-136
7x=56
x= 8 dozens of roses.
17-8 = 9 dozens of carnations
..
8*15+9*8=192

Answer by Alan3354(69443)

17 doz * $8 = $136.
For each $8 dozen replaced by a $15 dozen, the total price increases $7.
192 - 136 = 56
56/7 = 8
--> 9 of the $8, and 8 of the $15